How many pounds of chocolates costing $4.60 per pound may be mixed with 6 lb of chocolates costing $5 per pound to produce a mixture that can be sold for $4.90 per pound?
Answer by CharStar(110)

Chocolates costing - $4.60
Amount in lbs. - $4.60x
Total of Each - $4.60x
Chocolates costing - $5.00
Amount in lbs. - 6-x
Total of Each - 5.00(6-x)
Mixture - $4.90
Total of Mixutre - $4.90(6)
Equation
4.60x + 5.00(6-x) = 29.40
4.60x + 30.00 + 5.00x = 29.40
0.40x = 0.60
x = 1.5
Check your answer
4.60(1.5) = 6.9/It take 1.5 pounds of the chocolate costing $4.60
5.00(6-1.5) = 22.50/ It take 4.5 amounts of the chocolate costing $5.00
6.9 + 22.50 = 29.40
4.5 + 1.5 = 6.0
Hope that helps you